OpenStudy (anonymous):

If a curve with a radius of 88m is perfectly banked for a car traveling 75km/h, what must be the coefficient of static friction for a car not to skid when traveling at 95km/h? ac=v^2/r

OpenStudy (anonymous):

\[V=\sqrt{\mu*r*g}\] \[v^2 = \mu*r*g\] \[\mu=\frac{ v^2 }{ rg }\]

OpenStudy (anonymous):

Just Substitute and Find it out..)

OpenStudy (anonymous):